# Zero-shot Hierarchical Plant Segmentation via Foundation Segmentation Models and Text-to-image Attention

### v2

@@ 4.6 Ablation Studies
We use text-to-image cross-attention to determine keypoints, including the leaves’ bases. To evaluate the effectiveness of the cross-attention, we assess a straightforward alternative, where plant instances are grouped by clustering of the gravity centers of binary leaf masks instead of the base points. We use the same test dataset of PhenoBench [28] as the main experiment. The result is summarized in the top rows in Tab. 5. Using text-to-image cross-attention outperforms the method that only uses the gravity centers of clusters.
@@
To evaluate the effectiveness of the greedy clustering used for leaf grouping, we perform plant instance segmentation using the vanilla DBSCAN [18] instead of our greedy one. The bottom rows in Tab. 5 show that, although it is slightly inferior in $Prec_{50}$, our algorithm outperforms the usual DBSCAN in all other metrics.
@@
Table 6: Ablation study of using the GT leaf instances. We also list the results by our method using predicted leaf instances for reference. Using the GT instances indicates the upper-bound performance of our clustering-based plant individual segmentation.
@@ A.1 Cross-Attention and Weighted Least Square Line
Text-to-image cross-attention in one head is described as Eq. (5), where $\bm{F}_{I}\in\mathbb{R}^{N_{I}\times d}$ and $\bm{F}_{T}\in\mathbb{R}^{N_{T}\times d}$ denote the image feature and the text feature, respectively.
$\text{CrossAttention}(\bm{F}_{I},\bm{F}_{T})=\text{softmax}\left(\frac{\bm{F}_{I}\bm{F}_{T}^{\mathsf{T}}}{\sqrt{d}}\right).$~ (5)
In their official implementation  [^1] Grounding DINO, https://github.com/IDEA-Research/GroundingDINO, last accessed on July 10, 2025., the cross-attention module of Grounding DINO [16] consists of six cross-attention layers. We use the output of the final layer to obtain the feature map. Specifically, we extract four feature maps in resolution from $h/8\times w/8$ to $h/64\times w/64$ from each head, where $(h,w)$ denotes the size of the input image to the module. We resize to $h/8\times w/8$ and average them to obtain the feature map of the $i$-th leaf mask. We describe it as $f^{i}\in\mathbb{R}^{\frac{h}{8}\times\frac{w}{8}}_{+}$
@@
Now, we have the feature map of the $i$-th leaf mask, $f^{i}$. To determine the key points of the mask, we calculate the weighted least squares (WLS) line on the feature map and obtain key points as its edge points. Considering the value at the coordinate $(a,b)$ in the feature map $f^{i}[a,b]$ as the weight at the point $(a,b)$, an objective function of the WLS problem is described as Eq. (6), where $\bm{x}=[x_{1},x_{2}]^{\mathsf{T}}$ describes the WLS line.
@@
Solving Eq. (6) with $\frac{\partial J(\bm{x})}{\partial x_{1}}=\frac{\partial J(\bm{x})}{\partial x_{2}}=0$, the WLS line $\bm{x}$ can be calculated as Eq. (7). For more detailed explanations, please refer to [5].
$\bm{x}=(\bm{A}^{\mathsf{T}}\bm{W}\bm{A})^{-1}\bm{A}^{\mathsf{T}}\bm{W}\bm{b}.$ (7)
@@ A.2 Greedy Clustering
Figure 8: Comparison of ZeroPlantSeg and LeafOnlySAM. Since LeafOnlySAM does not involve plant segmentation, we only show the leaf instance segmentation results on the PhenoBench and Growliflower datasets. The left three columns show the results on the GrowliFlower dataset, and the right three columns show the results on the PhenoBench dataset.
Figure 9: Comparison of ZeroPlantSeg and Grounded SAM 2. The results are divided into two parts. The top rows show the comparison on the GrowliFlower dataset, and the bottom rows show the comparison on the PhenoBench dataset. Each picture is denoted by “Plant” or “Leaf” to indicate whether it is a leaf instance segmentation or a plant individual segmentation result.
Figure 10: Comparison of ZeroPlantSeg and Grounded SAM 2 on the SB20 dataset. Since SB20 only supplies annotations of individual plants, we only visualize the plant segmentation results.
@@ Appendix B Additional Results
While our method quantitatively surpasses these existing methods, Figs. 8,  9, and 10 show the visual results for the recent zero-shot models, namely, Grounded SAM2 [22] and LeafOnlySAM [31], compared to ours. Similar to the trend shown in the quantitative evaluation, our method more accurately estimates leaf and plant individual instances than the state-of-the-art zero-shot segmentation models.

### v3

@@ 4.6 Ablation Studies
We use text-to-image cross-attention to determine keypoints, including the leaves’ bases. To evaluate the effectiveness of the cross-attention, we assess a straightforward alternative, where plant instances are grouped by clustering of the gravity centers of binary leaf masks instead of the base points. We use the same test dataset of PhenoBench [PhenoBench] as the main experiment. The result is summarized in the top rows in Tab. 5. Using text-to-image cross-attention outperforms the method that only uses the gravity centers of clusters.
@@
To evaluate the effectiveness of the greedy clustering used for leaf grouping, we perform plant instance segmentation using the vanilla DBSCAN [DBSCAN] instead of our greedy one. The bottom rows in Tab. 5 show that, although it is slightly inferior in $Prec_{50}$, our algorithm outperforms the usual DBSCAN in all other metrics.
@@
Table 6: Ablation study of using the GT leaf instances. We also list the results of our method using predicted leaf instances for reference. Using the GT instances indicates the upper-bound performance of our clustering-based plant individual segmentation.
@@ Acknowledgments
[…]
Supplementary Material
This supplementary material provides additional information on technical details (Secs. A.1 and A.2). We also show additional visual results in Sec. B.
@@ A.1 Cross-Attention and Weighted Least Square Line
Text-to-image cross-attention in one head is described as Eq. (S5), where $\bm{F}_{I}\in\mathbb{R}^{N_{I}\times d}$ and $\bm{F}_{T}\in\mathbb{R}^{N_{T}\times d}$ denote the image feature and the text feature, respectively.
$\text{CrossAttention}(\bm{F}_{I},\bm{F}_{T})=\text{softmax}\left(\frac{\bm{F}_{I}\bm{F}_{T}^{\mathsf{T}}}{\sqrt{d}}\right).$~ (S5)
In their official implementation  [^1] Grounding DINO, https://github.com/IDEA-Research/GroundingDINO, last accessed on July 10, 2025., the cross-attention module of Grounding DINO [GDINO] consists of six cross-attention layers. We use the output of the final layer to obtain the feature map. Specifically, we extract four feature maps in resolution from $h/8\times w/8$ to $h/64\times w/64$ from each head, where $(h,w)$ denotes the size of the input image to the module. We resize to $h/8\times w/8$ and average them to obtain the feature map of the $i$-th leaf mask. We describe it as $f^{i}\in\mathbb{R}^{\frac{h}{8}\times\frac{w}{8}}_{+}$
@@
Now, we have the feature map of the $i$-th leaf mask, $f^{i}$. To determine the key points of the mask, we calculate the weighted least squares (WLS) line on the feature map and obtain key points as its edge points. Considering the value at the coordinate $(a,b)$ in the feature map $f^{i}[a,b]$ as the weight at the point $(a,b)$, an objective function of the WLS problem is described as Eq. (S6), where $\bm{x}=[x_{1},x_{2}]^{\mathsf{T}}$ describes the WLS line.
@@
Solving Eq. (S6) with $\frac{\partial J(\bm{x})}{\partial x_{1}}=\frac{\partial J(\bm{x})}{\partial x_{2}}=0$, the WLS line $\bm{x}$ can be calculated as Eq. (S7). For more detailed explanations, please refer to [wls].
$\bm{x}=(\bm{A}^{\mathsf{T}}\bm{W}\bm{A})^{-1}\bm{A}^{\mathsf{T}}\bm{W}\bm{b}.$ (S7)
@@ A.2 Greedy Clustering
Figure S8: Comparison of ZeroPlantSeg and LeafOnlySAM. Since LeafOnlySAM does not involve plant segmentation, we only show the leaf instance segmentation results on the PhenoBench and Growliflower datasets. The left three columns show the results on the GrowliFlower dataset, and the right three columns show the results on the PhenoBench dataset.
Figure S9: Comparison of ZeroPlantSeg and Grounded SAM 2. The results are divided into two parts. The top rows show the comparison on the GrowliFlower dataset, and the bottom rows show the comparison on the PhenoBench dataset. Each picture is denoted by “Plant” or “Leaf” to indicate whether it is a leaf instance segmentation or a plant individual segmentation result.
Figure S10: Comparison of ZeroPlantSeg and Grounded SAM 2 on the SB20 dataset. Since SB20 only supplies annotations of individual plants, we only visualize the plant segmentation results.
Figure S11: Failure cases of ZeroPlantSeg for leaf segmentation. The top row displays the incorrect split. The bottom row shows an example of failure to recognize leaves when they are occluded.
Figure S12: Failure cases of ZeroPlantSeg for plant segmentation. These show the situation where leaves from different plants are mistakenly seen as a whole.
@@ Appendix B Additional Results
While our method quantitatively surpasses these existing methods, Figs. S8,  S9, and S10 show the visual results for the recent zero-shot models, namely, Grounded SAM2 [ravi2024sam2segmentimages] and LeafOnlySAM [LSAM], compared to ours. Similar to the trend observed in the quantitative evaluation, our method more accurately estimates individual instances of leaves and plants than the state-of-the-art zero-shot segmentation models.
On the other hand, our method sometimes fails to recognize leaves where they are occluded by other leaves, or mistakenly splits one leaf into different parts (Fig. S12). This could be caused by the sliding windows strategy, as it would split one leaf into different windows and see them as individual leaves. In the meantime, it might also lead to incorrect plant segmentation by mixing leaves as if they were from the same plant. This could be caused by aggressive thresholds when executing Greedy DBSCAN (Fig. S12).
[…]